Calculus I, Fall 2013 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 11

1. Section 5.5: problems 35, 41, 43, 53, 57, 59, 67, 69, 79.

2. Section 5.6: problems 33, 43, 112, 117.